What is the theoretical probability of randomly selecting a 3 child family with 3 girls?
The gender of children is not a constant. It depends on the parental predispositions - mainly on the father's genes but apparently the mother's age can also affect the gender. The details of these influences are not understood well enough to develop an accurate model.
If you make the simplistic but unjustified assumption that both genders are equally likely, the answer is 1/8.
When all outcomes have the same probability of occurring the outcomes are?
"Equally likely"; nothing more, nothing less.
What is the probability of getting at least one head on three tosses of a coin?
In the experiment of tossing a fair coin 3 times, the sample space is made of the
following 8 equiprobable events: S = {HHH, THH, HTH, HHT, TTH, THT, HTT, TTT}.
The events that at least have one head are 7. So the probability of getting at least
one head on three tosses is: P(at least one H) = 7/8 = 0.875 = 87.5%
What is the probablity of picking a number less than 4 from a deck of cards?
With Aces high, the probability is 8/52 = 2/13.
What is the probablitiy of getting tails each time if a coin is tossed 4 times?
1/16 - if it is a fair coin.
What is the z score for 80 percent percent confidence?
80 percent percent = 0.008. The two tailed z-score for 0.008 is 2.652
What is the parameters that determine a Binomial Distribution?
The binomial distribution has two parameter, denoted by n and p.
n is the number of trials.
p is the constant probability of "success" at each trial.
How is the mean related to standard deviation?
Mean and standard deviation are not related in any way.
When two dice are rolled what is the probability of getting a sum of 9?
P(9) ≈ 11.11%
When two dice are rolled there are 36 equiprobable events.
The number of events that produce a sum of 9 are 4: (6,3), (5,4), (4,5) and (3,6).
So the probability of getting a sum of 9 is, P(9) = 4/36 = 1/9 = 0.1111.... ≈ 0.1111
≈ 11.11%
What is the probability of picking a heart or diamond from a pack of playing cards?
number of cards in a deck =52
number of cards that are heart =13
number of cards that are diamond =13
Probability that the card drawn is a heart or a diamond=13/52 +13/52 =26/52 or 1/2
Why two tail of normal distribution do not touch the horizontal axis?
The domain of the Normal distribution is the whole of the real line. As a result the horizontal axis is asymptotic to the Normal distribution curve. The curve gets closer and closer to the axis but never, ever reaches it.
What Z value corresponds to a confidences level of 93 percent?
Z = ± 1.8119 gives a two-tailed interval of 93%
How many permutations are in the word united?
UNITED = 6 letters
The letters in the word UNITED did not repeat so the
number of permutations = 6!
= 6x5x4x3x2
=720
How many even sums can you get from rolling two dice?
You can get 2, 4, 6, 8, 10, and 12, six even sums all together.
What is the probability of picking a jack and a diamond from a deck of cards?
1/52. Only one card, the Jack of diamonds, will satisfy your requirements.
Is it possible that in standard normal distribution standard deviation known and mean unknown?
it is possible to distribute standard deviation and mean but you dont have to understand how the mouse runs up the clock hicorky dickory dock.